# AutoAct: Automatic Agent Learning from Scratch via Self-Planning

### v2

@@ Appendix C Tool Library
  BingSearch
BingSearch engine can search for rich knowledge on the internet based on keywords, which can compensate for knowledge fallacy and knowledge outdated.
@@
BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query. Be specific and precise with your query to increase the chances of getting relevant results. For example, Bingsearch[popular dog breeds in the United States]
@@
Retrieve additional background knowledge crucial for tackling complex problems. It is especially beneficial for specialized domains like science and mathematics, providing context for the task
Retrieve[entity], which retrieves the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to retrieve. For example, Retrieve[Milhouse]
Lookup
A Lookup Tool returns the next sentence containing the target string in the page from the search tool, simulating Ctrl+F functionality on the browser.
Lookup[keyword], which returns the next sentence containing the keyword in the last passage successfully found by Retrieve or BingSearch. For example, Lookup[river].
Image2Text
Image2Text is used to detect words in images convert them into text by OCR and generate captions for images. It is particularly valuable when understanding an image semantically, like identifying objects and interactions in a scene.
Image2Text[image], which generates captions for the image and detects words in the image. You are recommended to use it first to get more information about the image to the question. If the question contains an image, it will return the caption and OCR text, else, it will return None. For example, Image2Text[image].
Text2Image
Text2Image Specializes in converting textual information into visual representations, facilitating the incorporation of textual data into image-based formats within the task.
Text2Image[text], which generates an image for the text provided by using multimodal models. For example, Text2Image[blue sky]
Code Interpreter
Code Interpreter is a tool or software that interprets and executes code written in Python. It analyzes the source code line by line and translates it into machine-readable instructions or directly executes the code and returns Execution results
Code[python], which interprets and executes Python code, providing a line-by-line analysis of the source code and translating it into machine-readable instructions. For instance, Code[print("hello world!")]
Table 5: Part of our tool library.
Appendix D Prompt
D.1 Prompt for Self-Instruct
See Table 6.
Prompt for Self-Instruct
I want you to be a QA pair generator to generate high-quality questions for use in Task described as follows:
Task Name: [task_name]
Task Description: [task_description]
Here are some Q&A pair examples from the Task:
[QA_pairs]
Modeled on all the information and examples above, I want you to generate new different [gen_num_per_round] Question-Answer pairs that cover a wide range of topics, some of which are difficult, some of which are easy, and require multiple steps of reasoning to get to the final answer. The format is like below:
[one_example]
Table 6: Prompt used for self-instruct.
D.2 Prompt for Tool Selection
See Table 7.
Prompt for Automatic Tool Selection
To successfully complete a complex task, the collaborative effort of three types of agents is typically required:
1. Plan Agent. This agent is used to plan the specific execution process of the benchmark, solving a given task by determining the order in which other expert language models are invoked;
2. Tool Agent. This agent is employed to decide how to use a specific tool when addressing a task. Tools encompass interactive tools within the task environment as well as external tools or models. The Tool Agent includes various tools that can be flexibly chosen;
3. Reflect Agent. This agent reflects on historical information and answers to assess whether the response aligns with the provided query.
Above all, the Tool Agent includes many tools that can be flexibly selected. Now your task is to select 3 tools from the Tool Library for solving a given task. Note that all tools are based on language models, and their inputs and outputs must be text. You only need to provide the names and descriptions of the tools in order, without any additional output.
Task Prompt Template
The following is the given task name and description, and you need to choose 3 corresponding tools from the Tool Library according to the above rules in the format of one line, one tool.
Task Name: [task_name]
Task Description: [task_description]
Tool Library: [list_of_tools]
Table 7: Prompt used for automatic tool selection.
D.3 Prompt for Trajectories Synthesis
See Table 8.
Prompt for Trajectories Synthesis
I expect you to excel as a proficient question answerer in the task.
Task Name: [task_name]
Task Description: [task_description]
Solve a question-answering task with interleaving Thought, Action, and Observation steps. Thought can reason about the current situation, and Action can be [action_num] types:
list of action selected from automatic tool selection [name, definition , usage]
Question: [question][scratchpad]
Table 8: Prompt used for trajectories synthesis.

### v4

@@ Appendix F Tool Library
  BingSearch BingSearch engine can search for rich knowledge on the internet based on keywords, which can compensate for knowledge fallacy and knowledge outdated. BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query. Be specific and precise with your query to increase the chances of getting relevant results. For example, Bingsearch[popular dog breeds in the United States]
Retrieve Retrieve additional background knowledge crucial for tackling complex problems. It is especially beneficial for specialized domains like science and mathematics, providing context for the task Retrieve[entity], which retrieves the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to retrieve. For example, Retrieve[Milhouse]
Lookup A Lookup Tool returns the next sentence containing the target string in the page from the search tool, simulating Ctrl+F functionality on the browser. Lookup[keyword], which returns the next sentence containing the keyword in the last passage successfully found by Retrieve or BingSearch. For example, Lookup[river].
Image2Text Image2Text is used to detect words in images convert them into text by OCR and generate captions for images. It is particularly valuable when understanding an image semantically, like identifying objects and interactions in a scene. Image2Text[image], which generates captions for the image and detects words in the image. You are recommended to use it first to get more information about the image to the question. If the question contains an image, it will return the caption and OCR text, else, it will return None. For example, Image2Text[image].
Text2Image Text2Image Specializes in converting textual information into visual representations, facilitating the incorporation of textual data into image-based formats within the task. Text2Image[text], which generates an image for the text provided by using multimodal models. For example, Text2Image[blue sky]
Code Interpreter Code Interpreter is a tool or software that interprets and executes code written in Python. It analyzes the source code line by line and translates it into machine-readable instructions or directly executes the code and returns Execution results Code[python], which interprets and executes Python code, providing a line-by-line analysis of the source code and translating it into machine-readable instructions. For instance, Code[print("hello world!")]
Table 6: Part of our tool library.
Appendix G Prompt
G.1 Prompt for Self-Instruct
See Tab. 7.
Prompt for Self-Instruct
I want you to be a QA pair generator to generate high-quality questions for use in Task described as follows: Task Name: [task_name] Task Description: [task_description] Here are some Q&A pair examples from the Task: [QA_pairs] Modeled on all the information and examples above, I want you to generate new different [gen_num_per_round] Question-Answer pairs that cover a wide range of topics, some of which are difficult, some of which are easy, and require multiple steps of reasoning to get to the final answer. The format is like below: [one_example]
Table 7: Prompt used for self-instruct.
G.2 Prompt for Tool Selection
See Tab. 8.
Prompt for Automatic Tool Selection
To successfully complete a complex task, the collaborative effort of three types of agents is typically required: 1. Plan Agent. This agent is used to plan the specific execution process of the benchmark, solving a given task by determining the order in which other expert language models are invoked; 2. Tool Agent. This agent is employed to decide how to use a specific tool when addressing a task. Tools encompass interactive tools within the task environment as well as external tools or models. The Tool Agent includes various tools that can be flexibly chosen; 3. Reflect Agent. This agent reflects on historical information and answers to assess whether the response aligns with the provided query. Above all, the Tool Agent includes many tools that can be flexibly selected. Now your task is to select 3 tools from the Tool Library for solving a given task. Note that all tools are based on language models, and their inputs and outputs must be text. You only need to provide the names and descriptions of the tools in order, without any additional output.
Task Prompt Template
The following is the given task name and description, and you need to choose 3 corresponding tools from the Tool Library according to the above rules in the format of one line, one tool. Task Name: [task_name] Task Description: [task_description] Tool Library: [list_of_tools]
Table 8: Prompt used for automatic tool selection.
G.3 Prompt for Trajectories Synthesis
See Tab. 9.
Prompt for Trajectories Synthesis
I expect you to excel as a proficient question answerer in the task. Task Name: [task_name] Task Description: [task_description] Solve a question-answering task with interleaving Thought, Action, and Observation steps. Thought can reason about the current situation, and Action can be [action_num] types: list of action selected from automatic tool selection [name, definition , usage] Question: [question][scratchpad]
Table 9: Prompt used for trajectories synthesis.
Appendix H Database Cases
HotpotQA:
Question: The deepest part of the ocean, is located in which ocean?
Answer: The Pacific Ocean
Question: The famous scientist who discovered gravity, lived in which century?
Answer: 17th century
Question: The first successful flight of a power was made by which inventor?
Answer: The Wright brothers
Question: The highest mountain peak in the solar system is located on which planet?
Answer: Mars
Question: In the novel "Pride and Prejudice", what is the name of Mr. Darcy’s estate in Derbyshire, England?
Answer: Pemberley
ScienceQA:
Question: Which of the following is a type of renewable energy?
Options: (A) Coal (B) Oil (C) Natural gas (D) Solar power
Caption: A picture of a solar cell
Answer: D. Solar power
Question: Which of the following is the term for the process by which the Earth’s weather patterns are influenced by the movement of air in the atmosphere?
Options: (A) Weathering (B) Erosion (C) Deposition (D) Atmospheric circulation
Caption: An image of air currents in the atmosphere
Answer: D. Atmospheric circulation
Question: Which of the following is a type of chemical reaction that involves the transfer of electrons between atoms?
Options: (A) Combustion (B) Photosynthesis (C) Respiration (D) Electrolysis
Caption: An image of a battery
Answer: D. Electrolysis
Question: Which of the following is an example of a type of weather phenomenon that occurs when warm air rises and cool air sinks?
Options: (A) Thunderstorms (B) Hurricanes (C) Fog (D) Fronts
Caption: An image of a front
Answer": D. Fronts
Question: Which of the following is the term for the process by which water is purified through the use of microorganisms that consume organic matter?
Options: (A) Filtration (B) Sedimentation (C) Biodegradation (D) Disinfection
Caption: An image of a water treatment plant
Answer: C. Biodegradation
Appendix I Training Data Example
Here we give an example of the training data for each sub-agent.
Plan-Agent (generate Thought):
Input:
(format requirements) (tool usage instructions)
Question: The first human-made object to land on the moon, in 1969, was which spacecraft?
Thought: I should first search the Moon landing history.
Action: BingSearch[moon landing spacecraft]
Observation: A Moon landing or lunar landing is the arrival of a spacecraft on the surface of the Moon. The first human-made object to touch the Moon was the Soviet Union’s Luna 2, on 13 September 1959. Apollo 11 (July 16–24, 1969) was the American spaceflight that ……
Thought:
Output:
Retrieve for more information about Apollo 11
Plan-Agent (generate Action name):
Input:
(format requirements) (tool usage instructions)
Question: The first human-made object to land on the moon, in 1969, was which spacecraft?
Thought: I should first search the Moon landing history.
Action: BingSearch[moon landing spacecraft]
Observation: A Moon landing or lunar landing is the arrival of a spacecraft on the surface of the Moon. The first human-made object to touch the Moon was the Soviet Union’s Luna 2, on 13 September 1959. Apollo 11 (July 16–24, 1969) was the American spaceflight that ……
Thought: Retrieve for more information about Apollo 11.
Action:
Output:
@@ Appendix I Training Data Example
Tool-Agent:
Input:
(format requirements) (tool usage instructions)
Question: The first human-made object to land on the moon, in 1969, was which spacecraft?
Thought: I should first search the Moon landing history.
Action: BingSearch[moon landing spacecraft]
Observation: A Moon landing or lunar landing is the arrival of a spacecraft on the surface of the Moon. The first human-made object to touch the Moon was the Soviet Union’s Luna 2, on 13 September 1959. Apollo 11 (July 16–24, 1969) was the American spaceflight that ……
Thought: Retrieve for more information about Apollo 11.
Action: Retrieve
Output:
Apollo 11
Reflect-Agent (generate Thought):
Input:
(format requirements) (tool usage instructions)
Question: The first human-made object to land on the moon, in 1969, was which spacecraft?
Thought: I should first search the Moon landing history.
Action: BingSearch[moon landing spacecraft]
Observation: A Moon landing or lunar landing is the arrival of a spacecraft on the surface of the Moon. The first human-made object to touch the Moon was the Soviet Union’s Luna 2, on 13 September 1959. Apollo 11 (July 16–24, 1969) was the American spaceflight that ……
Thought: Retrieve for more information about Apollo 11.
Action: Retrieve[Apollo 11]
Action: Finish[Apollo 11]
Observation: Please reflect your answer based on the history.
Thought:
Output:
The question asks about the first human-made object to land on the moon, so it seems that the Soviet Union’s Luna 2 is more like the answer.
Reflect-Agent (generate Action):
Input:
(format requirements) (tool usage instructions)
Question: The first human-made object to land on the moon, in 1969, was which spacecraft?
Thought: I should first search the Moon landing history.
Action: BingSearch[moon landing spacecraft]
Observation: A Moon landing or lunar landing is the arrival of a spacecraft on the surface of the Moon. The first human-made object to touch the Moon was the Soviet Union’s Luna 2, on 13 September 1959. Apollo 11 (July 16–24, 1969) was the American spaceflight that ……
Thought: Retrieve for more information about Apollo 11.
Action: Retrieve[Apollo 11]
Action: Finish[Apollo 11]
Observation: Please reflect your answer based on the history.
Thought: The question asks about the first human-made object to land on the moon, so it seems that the Soviet Union’s Luna 2 is more like the answer.
Action:
Output:
Reflect[wrong]